OpenStudy (anonymous):

Find the integral of (4-x^2)^(-3/2) dx from 0 to 1. (Answer: sqrt(3)/12) (I got [-2(4-x^2)^(-1/2)] from 0 to 1 but that's not the answer.)

OpenStudy (anonymous):

Is there a easier way to do this?

OpenStudy (abb0t):

You could of used Wolfram. That's easier method.

OpenStudy (anonymous):

I thought I did change the limits @Loser66... you could change the bottom one to 0. @Idealist, have you learned about trig substitution yet?

OpenStudy (anonymous):

I think I did.

OpenStudy (anonymous):

Let me try one more time... lol

OpenStudy (abb0t):

OHHHH! He said your method is unfamiliar and implied that your method was dumb!! Buuuuuuuuurn!!!

OpenStudy (anonymous):

\[\int_0^1\left(4-x^2\right)^{-3/2}~dx\] This integral is with respect to \(x\), which changes from 0 to 1. If we let \(x=2\sin u\), then \(u=\arcsin \dfrac{x}{2}\), which changes from \(\arcsin 0=0\) to \(\arcsin \dfrac{1}{2}=\dfrac{\pi}{6}\). \[\int_0^{\pi/6}(4-4\sin^2u)^{-3/2}~2\cos u~du\]

OpenStudy (anonymous):

So, @Idealist, the procedure here is to reduce to a more easily integrable expression: \[\frac{1}{4}\int_0^{\pi/6}(1-\sin^2u)^{-3/2}~\cos u~du\\ \frac{1}{4}\int_0^{\pi/6}(\cos^2u)^{-3/2}~\cos u~du\\ \frac{1}{4}\int_0^{\pi/6}\sec^2u~du\\ \]

OpenStudy (dan815):

lmao

OpenStudy (dan815):

abbot

OpenStudy (anonymous):

And then you integrate, right?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Wow, I like this method more.

OpenStudy (loser66):

@SithsAndGiggles Thanks for giving us the easiest way to get the answer.

OpenStudy (anonymous):

This is the hardest integral that I've seen!

OpenStudy (dan815):

hahaa

OpenStudy (anonymous):

Don't laugh.

OpenStudy (dan815):

<3

OpenStudy (loser66):

@Idealist sweetie, ignore him

OpenStudy (raffle_snaffle):

Hahaha